# Supporting

### v1

@@ 2.2 Sequential classification learning for quantile OTRs with continuous outcome
Building on the preceding discussion, the estimation of quantile OTRs $d_{q^{*}}^{*}$ can be carried out by first estimating the optimal quantile $q^{*}$, and then using this value to construct an estimate of $d_{q^{*}}^{*}$. The estimation of $q^{*}$ can be obtained via the previously described binary search procedure, wherein the optimal survival functions $\max_{d}S(q,d)$ are replaced by their empirical estimates. Since the survival function can be expressed as an expectation $S(q,d)=E[I\{Y(d)>q\}]$, we can borrow the idea of classification learning (Zhang et al., 2012a, ) to estimate $\max_{d}S(q,d)$ and ultimately $d_{q^{*}}^{*}=\arg\max_{d}S(q^{*},d)$, taking advantage of the favorable optimization and statistical properties inherent to classification learning. Overall, the estimation of quantile OTRs is effectively recast as a sequential classification learning problem.
@@
A classification learning framework with the doubly robust property is adopted for $d_{q}^{*}$ (Zhang et al., 2012a, ), which is underpinned by the identification of $S(q,d)$ Robins et al., (1994):
@@
The doubly robust classification idea is to reformulate the ${d}_{q}^{*}$ as a solution to the weighted classification problem Zhang et al., 2012a :
${d}_{q}^{*}=$ $\operatorname*{arg\,max}_{d}{S}(q,d)$
$=$ $\operatorname*{arg\,min}_{d}E[|{C}(\boldsymbol{X},q;{\pi}^{*},{g}^{*})|I\{Z^{*}\neq d(\boldsymbol{X})\}],$ (4)
@@
Note that regimes $d(\boldsymbol{X})$ can always be represented as $I\{f(\boldsymbol{X})>0\}$ for some decision functions $f$. To facilitate the optimization, the classification approach proceeds by utilizing the convex hinge loss $\phi\{(2{Z}^{*}-1)f(\boldsymbol{X})\}$ to replace the non-convex 0-1 loss $I\{Z^{*}\neq d(\boldsymbol{X})\}$ Zhao et al., (2012), where the hinge loss function $\phi(x)=(1-x)^{+}=\max\{1-x,0\},x\in\mathbb{R}$. Then, we can estimate $d_{q}^{*}(\boldsymbol{x})$ by $\hat{d}_{q}(\boldsymbol{x})=I\{\hat{f}_{q}(\boldsymbol{x})>0\}$, where
@@
Here, $\hat{\pi}$ is an estimator for the propensity score $\pi^{*}$, and $\hat{g}$ is an estimator for the conditional survival function $g^{*}$, where the estimation of these nuisance functions is discussed in Section 2.4. Additionally, $\widehat{Z}_{i}=I\{{C}(\boldsymbol{X}_{i},q;\hat{\pi},\hat{g})>0\}$ is the empirical analogue of $Z_{i}^{*}$, constructed by replacing the nuisance functions by their estimators. The regularity term $\lambda_{n}\|f\|^{2}$ is added to avoid overfitting with $\lambda_{n}$ as the tuning parameter. We also consider a prespecified set of decision functions =$\mathcal{F}=\{b+h(\boldsymbol{x}):b\in\mathbb{R},h(\boldsymbol{x})\in\mathcal{H}_{k}\}$ and the norm $\|f\|^{2}=\|h\|_{k}^{2}$, where $\mathcal{H}_{k}$ denotes a reproducing kernel Hilbert space (RKHS) and $\|h\|_{k}$ is the norm induced by the inner product of the RKHS. To derive linear quantile OTRs, we consider $\mathcal{H}_{k}$ as the RKHS reproduced by a linear kernel $\left\langle\boldsymbol{x},\boldsymbol{z}\right\rangle$, where $\boldsymbol{x},\boldsymbol{z}\in\mathcal{X}$. For nonlinear quantile OTRs, we adopt the Gaussian kernel, $k_{\sigma_{n}}(\boldsymbol{x},\boldsymbol{z})=\exp\left(-\sigma_{n}^{2}\|\boldsymbol{x}-\boldsymbol{z}\|^{2}\right),$ where $\boldsymbol{x},\boldsymbol{z}\in\mathcal{X}$ and the parameter $\sigma_{n}$ controls the complexity of the function reproduced by the kernel. The parameters $\lambda_{n}$ (and $\sigma_{n}$) can be chosen via $5$-fold cross-validation. For further details on solving the minimization problem (5), we refer the reader to Zhao et al., (2012).
Building on the identification (3), we adopt the following estimator for $\max_{d}S(q,d)=S(q,d_{q}^{*})$ Jiang et al., (2017):
@@
where $\Phi(x),x\in\mathbb{R},$ denotes the cumulative distribution function of the standard normal distribution, and $h_{n}$ serves as a bandwidth parameter. This estimator employs a soft estimated regime $\Phi\{\hat{f}_{q}(\boldsymbol{x})/h_{n}\}$ rather than the hard estimated regime $\hat{d}_{q}(\boldsymbol{x})=I\{\hat{f}_{q}(\boldsymbol{x})>0\}$, promoting a more stable evaluation of $\max_{d}S(q,d)$. For the choice of the bandwidth parameter $h_{n}$, we recommend $h_{n}=0.2/\log n$, which performs well in our simulation and satisfies the conditions to establish the consistency result in Theorem 3.1.
@@
We have summarized our sequential classification learning procedure in Algorithm 1. There are several tuning parameters $l^{1}$, $r^{1}$, $\kappa_{n}$, and $\epsilon_{n}$ that need to be properly selected for the binary search procedure. The $l^{1}$ and $r^{1}$ should ensure that $q^{*}$ is included in the interval $(l^{1},r^{1})$. If the support of $Y$ is a bounded interval, we can choose $(l^{1},r^{1})$ that encompasses the support of $Y$. Otherwise, we can empirically choose $l^{1}$ and $r^{1}$ as the minimum and the maximum of the observed outcome, respectively. The $\kappa_{n}$ and $\epsilon_{n}$ should be small to avoid introducing a large bias when estimating the optimal quantile $q^{*}$. We recommend $\kappa_{n}=6^{-1}n^{-1/2}\operatorname{sd}(Y)$ and $\epsilon_{n}=0.5n^{-1/2}$, where $\operatorname{sd}(Y)$ is the sample standard deviation of random variable $Y$. It performs well in our simulation and satisfies the conditions to establish the consistent result in Theorem 3.1.
@@
2) For $b=1,2,...,$ repeat steps (2.1)-(2.6) until $r^{b}-l^{b}\leq\kappa_{n}$.
@@
 (2.2) For a series of tuning parameters $\lambda_{n}$ (and $\sigma_{n}$ if we choose a Gaussian kernel), choose the best parameter via 5-fold cross-validation.
@@
2.3 Sequential classification learning for discrete outcomes
@@
$S^{m}(q,d)=\left\{\begin{array}[]{ll}1&\text{ if }q\leq v_{0},\\ S(v_{k},d)&\text{ if }q=v_{k},k=1,2,\ldots,l-1,\\ \lambda S(v_{k},d)+(1-\lambda)S(v_{k+1},d)&\text{ if }q=\lambda v_{k}+(1-\lambda)v_{k+1},0<\lambda<1,\\ &k=0,1,\ldots,l-1,\\ 0&\text{ if }q\geq v_{l}.\end{array}\right.$ (7)
@@
The key idea of our approach is to map the discrete outcome to a continuous outcome by uniformly distributing the mass at $v_{j}$ over the interval $[v_{j-1},v_{j}]$. Define the set of regimes maximizing the smoothed quantile by $\tilde{\mathcal{D}}^{*}=\{\tilde{d}:Q_{\tau}^{m}\{Y(\tilde{d})\}=\max_{d}Q_{\tau}^{m}\{Y(d)\}\}$ and recall the set of quantile OTRs $\mathcal{D}^{*}=\{\tilde{d}:Q_{\tau}\{Y(\tilde{d})\}=\max_{d}Q_{\tau}\{Y(d)\}\}$. Surprisingly, we can prove that $\tilde{\mathcal{D}}^{*}\subseteq\mathcal{D}^{*}$ (see Theorem LABEL:the:smoothed in Section LABEL:s6.6 of the supplementary material). Therefore, the regime $\tilde{d}^{*}\in\tilde{\mathcal{D}}^{*}$ belongs to the solution set $\mathcal{D}^{*}$. Then, since $S^{m}(q,d)$ is a continuous function of $q$, we can consistently estimate the quantile OTRs using Algorithm 1, substituting the estimated optimal decision functions $\hat{f}_{q}$ and the estimated optimal survival functions $\widehat{S}_{h_{n}}(q,\hat{f}_{q})$ with their smoothed counterparts, $\hat{f}_{q}^{m}$ and $\widehat{S}_{h_{n}}^{m}(q,\hat{f}_{q}^{m})$, obtained through survival smoothing.
@@
where $\psi^{m}(a,q;{\pi}^{*},{g}^{*})$ is the smoothed counterpart of $\psi(a,q;{\pi}^{*},{g}^{*})$. The explicit expression for $\psi^{m}(a,q;{\pi}^{*},{g}^{*})$ is given in (LABEL:eq:psim) of the supplementary material. Then, we apply the classification idea to derive $\hat{f}_{q}^{m}$ and subsequently construct $\widehat{S}_{h_{n}}^{m}(q,\hat{f}_{q}^{m})$ following a similar derivation in Section 2.2. We have provided detailed procedures to compute $\hat{f}_{q}^{m}$ and $\widehat{S}_{h_{n}}^{m}(q,\hat{f}_{q}^{m})$ in Section LABEL:s7 of the supplementary material. Additionally, a detailed algorithm tailored for discrete outcomes is presented in Algorithm LABEL:alg_2 of the supplementary material.
A key advantage of focusing on $\tilde{\mathcal{D}}^{*}$ from the solution set $\mathcal{D}^{*}$ is that it can exclude quantile OTRs that are close to the non-optimal regime and therefore improve the performance when estimating quantile OTRs. When the homoscedastic treatment effect assumption (Assumption 9 in Section 3.2) holds, only the effective quantile OTRs remain in the set $\tilde{\mathcal{D}}^{*}$ (see Theorem LABEL:the:expectation in Section LABEL:s6.6 of the supplementary material), which helps solve the ineffectiveness issue.
@@ 3.1 Result for continuous outcomes
Assumption 2: The counterfactual survival functions under treatment $1$ and 00, $S({q},1)$ and $S({q},0)$, are continuous functions of $q$.
Assumption 3: The nuisance models $\hat{\pi}(a|\boldsymbol{x}),\hat{g}(q;\boldsymbol{x},a)$ as functions of $q$, $\boldsymbol{x}$, and $a$ are in the Vapnik–Chervonenkis (VC) classes $\mathcal{F}_{\pi}$ and $\mathcal{F}_{g}$, respectively.
@@
Assumption 5: For each fixed $(\boldsymbol{x},a)$, $\hat{\pi}(a|\boldsymbol{x})$ and $\hat{g}(q;\boldsymbol{x},a)$ converge in probability to $\pi^{0}(a|\boldsymbol{x})$ and $g^{0}(q;\boldsymbol{x},a)$ uniformly in $q\in(l,r)$, where the interval $(l,r)$ contains $(l^{1},r^{1})$ as defined in Algorithm 1. Moreover, there exists a constant $\gamma>0$ such that,
@@
These assumptions were used in the related literature when a classification method is employed to estimate the OTRs Zhao et al., 2015b . VC class (Assumption 3) includes regular parametric classes (e.g., the conditional Gaussian model and the generalized linear model) and infinite-dimensional classes when certain smoothness or boundedness are satisfied Van Der Vaart and Wellner, (1996). Assumption 4 is required to ensure the estimator is well-defined with nonzero denominators. Assumption 5 implies that $\hat{\pi}$ and $\hat{g}$ converge to fixed functions $\pi^{0}$ and $g^{0}$, and $\pi^{0}$ and $g^{0}$ are not required to match the true functions $\pi^{*}$ and $g^{*}$. If we utilize parametric models to estimate $\pi^{*}$ and $g^{*}$, then $\gamma=1/2$ in Assumption 5.
@@
$\max\{-C(\boldsymbol{X},q;{\pi}^{0},{g}^{0}),0\}$, and $C(\boldsymbol{X},q;{\pi}^{0},{g}^{0})$ is defined in (4). When either ${\pi}^{0}=\pi^{*}$ or ${g}^{0}=g^{*}$, $E\{{C}_{+}(\boldsymbol{X},q)|\boldsymbol{X}=\boldsymbol{x},A=1\}-E\{{C}_{-}(\boldsymbol{X},q)|\boldsymbol{X}=\boldsymbol{x},A=0\}=\tau(\boldsymbol{x},q)$. Furthermore, we introduce the sets $\mathcal{X}^{+}_{q}=\{\boldsymbol{x}\in\mathcal{X}:2\eta(\boldsymbol{x},q)-1>0\}$ and $\mathcal{X}^{-}_{q}=\{\boldsymbol{x}\in\mathcal{X}:2\eta(\boldsymbol{x},q)-1<0\}$. A distance function to the boundary between $\mathcal{X}^{+}_{q}$ and $\mathcal{X}^{-}_{q}$ is denoted by $\Delta(\boldsymbol{x},q)=\tilde{d}\left(\boldsymbol{x},\mathcal{X}^{+}_{q}\right)$ if $\boldsymbol{x}$ belongs to $\mathcal{X}^{-}_{q}$, $\Delta(\boldsymbol{x},q)=\tilde{d}\left(\boldsymbol{x},\mathcal{X}^{-}_{q}\right)$ if $\boldsymbol{x}$ lies within $\mathcal{X}^{+}_{q}$, and $\Delta(\boldsymbol{x},q)=0$ otherwise, where $\tilde{d}(\boldsymbol{x},\mathcal{O})$ represents the Euclidean norm-based distance of $\boldsymbol{x}$ to set $\mathcal{O}$. Recall that ${q}^{*}=\max_{d}Q_{\tau}\{Y({d})\}$ and $d_{q}^{*}=\operatorname*{arg\,max}_{d}S(q,d)$. Define, for $\delta>0$, $Q_{\delta}=\{q:|S(q,d_{q}^{*})-1+\tau|<\delta\}.$
@@
In the related literature, assumptions similar to Assumptions 6–8 are typically stated at a fixed value $q$ Zhao et al., (2012); Zhou et al., (2017); Shi et al., (2020); Ma et al., (2011). We generalize these assumptions to hold over a neighborhood $Q_{\delta^{*}}$. Since $\delta^{*}$ can be very small, our assumptions are only slightly stronger than those in the existing literature. We require such stronger assumptions to accommodate that the final classification to derive the estimated quantile OTR $\hat{d}_{\hat{q}}$ relies on the estimated optimal quantile $\hat{q}$ rather than a fixed $q$.
Assumption 6 is introduced to guarantee the proper behavior of the classification procedure deriving $\hat{f}_{q}$ in Algorithm 1. The geometric noise exponent $\beta$ (Assumption 6) describes how fast the density of the distance $\Delta(\boldsymbol{X},{q}^{*})$ decays near the decision boundary, i.e., the $\boldsymbol{x}$ satisfying $02\eta(\boldsymbol{x},{q}^{*})-1=$ 0. When the decay is rapid, there are fewer data points near the boundary, making classification easier, and allowing for the selection of a larger geometric noise exponent $\beta$. An example is distinctly separable data, that is, when $|2\eta(\boldsymbol{X},{q}^{*})-1|>\delta_{0}>0$ with probability $1$ for some constant $\delta_{0}$, and $\eta(\boldsymbol{x},{q})$ is continuous, $\beta$ can be arbitrarily large. Assumption 7 is introduced to ensure the evaluation procedure in Algorithm 1 works properly, i.e., to bound the difference between $\widehat{S}_{h_{n}}(q,\hat{f}_{q})$ and $S(q,d_{q}^{*})$. When $\tau(\boldsymbol{X},{q}^{*})$ has a uniformly bounded density function near 00, and $\tau(\boldsymbol{x},{q})$ is continuous, Assumption 7 holds with $\eta=1$. If there exists some $\delta_{0}>0$ such that $|\tau(\boldsymbol{X},{q}^{*})|>\delta_{0}$ with probability $1$, and $\tau(\boldsymbol{x},{q})$ is continuous, then Assumption 7 holds with an arbitrary large $\eta$. In Assumption 8, for a given $q$, when $S(q_{0},d_{q}^{*})$ is differentiable at point $q_{0}={q}^{*}$ with a nonzero derivative, i.e., the density function for $Y(d_{q}^{*})$ at ${q}^{*}$ is nonzero, both upper and lower bounds for $|S(q_{0},d_{q}^{*})-S({q}^{*},d_{q}^{*})|$ in Assumption 8 can be established with a proper choice of $\delta^{*}$. Thus, Assumption 8 imposes a weaker condition than requiring differentiability.
For simplicity in notations, we also assume $\lambda_{n}$, $\sigma_{n}$, and $h_{n}$ are the same across different ${q}$.

### v2

@@ 2.2 Successive classification learning for continuous outcomes
Building on the preceding discussion, the estimation of quantile OTRs $d_{q^{*}}^{*}$ can be carried out by first estimating the optimal quantile $q^{*}$, and then using this value to construct an estimate of $d_{q^{*}}^{*}$. The estimation of $q^{*}$ can be obtained via the previously described binary search procedure, wherein the optimal survival functions $\max_{d}S(q,d)$ are replaced by their empirical estimates. Since the survival function can be expressed as an expectation $S(q,d)=E[I\{Y(d)>q\}]$, we can borrow the idea of classification learning (zhang2012estimating) to estimate $\max_{d}S(q,d)$ and ultimately $d_{q^{*}}^{*}=\arg\max_{d}S(q^{*},d)$, taking advantage of the favorable optimization and statistical properties inherent to classification learning. Overall, the estimation of quantile OTRs is effectively recast as a successive classification learning problem.
@@
A classification learning framework with the doubly robust property is adopted for $d_{q}^{*}$ (zhang2012estimating), which is underpinned by the identification of $S(q,d)$ robins1994estimation:
@@
The doubly robust classification idea is to reformulate the ${d}_{q}^{*}$ as a solution to the weighted classification problem zhang2012estimating:
${d}_{q}^{*}=\operatorname*{arg\,max}_{d}{S}(q,d)=\operatorname*{arg\,min}_{d}E[|{C}(\boldsymbol{X},q;{\pi}^{*},{g}^{*})|I\{Z^{*}\neq d(\boldsymbol{X})\}],$ (4)
@@
Note that regimes $d(\boldsymbol{X})$ can always be represented as $I\{f(\boldsymbol{X})>0\}$ for some decision function $f$. The classification approach then employs the convex hinge loss $\phi\{(2Z^{*}-1)f(\boldsymbol{X})\}$ as a surrogate for the non-convex 0-1 loss $I\{Z^{*}\neq d(\boldsymbol{X})\}$ zhao2012estimating, where the hinge loss function is defined as $\phi(x)=(1-x)_{+}=\max\{1-x,0\},\,x\in\mathbb{R}$. Replacing the 0-1 loss with the hinge loss offers several advantages. First, the convexity of the hinge loss enables efficient optimization. Second, the hinge loss $\phi\{(2Z^{*}-1)f(\boldsymbol{X})\}=\{1-(2Z^{*}-1)f(\boldsymbol{X})\}_{+}$ remains positive not only when $\boldsymbol{X}$ is misclassified, but also when $\boldsymbol{X}$ is correctly classified yet lies close to the decision boundary cortes1995support. This property encourages the samples $\{\boldsymbol{X}_{i}\}_{i=1}^{n}$ to be pushed farther away from the decision boundary, thereby enhancing both robustness and generalization of the estimated regime.
With the hinge loss function, we can estimate $d_{q}^{*}(\boldsymbol{x})$ by $\hat{d}_{q}(\boldsymbol{x})=I\{\hat{f}_{q}(\boldsymbol{x})>0\}$, where
@@
Here, $\hat{\pi}$ is an estimator for the propensity score $\pi^{*}$, and $\hat{g}$ is an estimator for the conditional survival function $g^{*}$, where the estimation of these nuisance functions is discussed in Section 2.4. Additionally, $\widehat{Z}_{i}=I\{{C}(\boldsymbol{X}_{i},q;\hat{\pi},\hat{g})>0\}$ is the empirical analogue of $Z_{i}^{*}$, constructed by replacing the nuisance functions by their estimators. The regularity term $\lambda_{n}\|f\|^{2}$ is added to avoid overfitting with $\lambda_{n}$ as the tuning parameter. We also consider a prespecified set of decision functions $\mathcal{F}=\{b+h(\boldsymbol{x}):b\in\mathbb{R},h(\boldsymbol{x})\in\mathcal{H}_{k}\}$ and the norm $\|f\|^{2}=\|h\|_{k}^{2}$, where $\mathcal{H}_{k}$ denotes a reproducing kernel Hilbert space (RKHS) and $\|h\|_{k}$ is the norm induced by the inner product of the RKHS. To derive linear quantile OTRs, we consider $\mathcal{H}_{k}$ as the RKHS reproduced by a linear kernel $\left\langle\boldsymbol{x},\boldsymbol{z}\right\rangle$, where $\boldsymbol{x},\boldsymbol{z}\in\mathcal{X}$. For nonlinear quantile OTRs, we adopt the Gaussian kernel, $k_{\sigma_{n}}(\boldsymbol{x},\boldsymbol{z})=\exp\left(-\sigma_{n}^{2}\|\boldsymbol{x}-\boldsymbol{z}\|^{2}\right),$ where $\boldsymbol{x},\boldsymbol{z}\in\mathcal{X}$ and the parameter $\sigma_{n}$ controls the complexity of the function reproduced by the kernel. The parameters $\lambda_{n}$ (and $\sigma_{n}$) can be chosen via 10-fold cross-validation. For further details on solving the minimization problem (5), we refer the reader to zhao2012estimating.
Building on the identification (3), we adopt the following estimator for $\max_{d}S(q,d)=S(q,d_{q}^{*})$ jiangDoublyRobustEstimation2017:
@@
where $\Phi(x),x\in\mathbb{R},$ denotes the cumulative distribution function of the standard normal distribution, and $h_{n}$ serves as a bandwidth parameter. This estimator employs a soft estimated regime $\Phi\{\hat{f}_{q}(\boldsymbol{x})/h_{n}\}$ rather than the hard estimated regime $\hat{d}_{q}(\boldsymbol{x})=I\{\hat{f}_{q}(\boldsymbol{x})>0\}$ to facilitate the theoretical development. For the choice of the bandwidth parameter $h_{n}$, we recommend $h_{n}=0.2/\log n$, which performs well in our simulation and satisfies the conditions to establish the consistency result in Theorem 3.1.
Remark 2.1.
The soft regime $\Phi\{\hat{f}_{q}(\boldsymbol{x})/h_{n}\}$ to construct $\widehat{S}_{h_{n}}(q,\hat{f}_{q})$ in (6) may be replaced by the hard regime $I\{\hat{f}_{q}(\boldsymbol{x})>0\}$, although no theoretical guarantees exist as the indicator function $I\{\hat{f}_{q}(\boldsymbol{x})>0\}$ possesses an inherent jump discontinuity. A simulation provided in the Section S3.4 of the supplementary material reveals that these two regimes exhibit similar performance. Thus, we favor the soft regime for its theoretical guarantee.
@@
We have summarized our successive classification learning procedure in Algorithm 1. There are several tuning parameters $l^{1}$, $r^{1}$, $\kappa_{n}$, and $\epsilon_{n}$ that need to be properly selected for the binary search procedure. The $l^{1}$ and $r^{1}$ should ensure that $q^{*}$ is included in the interval $(l^{1},r^{1})$. If the support of $Y$ is a bounded interval, we can choose $(l^{1},r^{1})$ that encompasses the support of $Y$. Otherwise, we can empirically choose $l^{1}$ and $r^{1}$ as the minimum and the maximum of the observed outcome, respectively. The $\kappa_{n}$ and $\epsilon_{n}$ should be small to avoid introducing a large bias when estimating the optimal quantile $q^{*}$. We recommend $\kappa_{n}=6^{-1}n^{-1/2}\operatorname{sd}(Y)$ and $\epsilon_{n}=0.5n^{-1/2}$, where $\operatorname{sd}(Y)$ is the sample standard deviation of random variable $Y$. It performs well in our simulation and satisfies the conditions to establish the consistent result in Theorem 3.1.
@@
2) For b=1,2,,b=1,2,..., repeat steps (2.1)-(2.6) until $r^{b}-l^{b}\leq\kappa_{n}$.
@@
 (2.2) For a series of tuning parameters $\lambda_{n}$ (and $\sigma_{n}$ if we choose a Gaussian kernel), choose the best parameter via 10-fold cross-validation.
@@
2.3 Successive classification learning for discrete outcomes
@@
[…]
@@
The key idea of our approach is to map the discrete outcome to a continuous outcome by uniformly distributing the mass at $v_{j}$ over the interval $v_{j-1},v_{j}]$. Define the set of regimes maximizing the smoothed quantile by $\tilde{\mathcal{D}}^{*}=\{\tilde{d}:Q_{\tau}^{m}\{Y(\tilde{d})\}=\max_{d}Q_{\tau}^{m}\{Y(d)\}\}$ and recall the set of quantile OTRs $\mathcal{D}^{*}=\{\tilde{d}:Q_{\tau}\{Y(\tilde{d})\}=\max_{d}Q_{\tau}\{Y(d)\}\}$. Surprisingly, we can prove that $\tilde{\mathcal{D}}^{*}\subseteq\mathcal{D}^{*}$ (see Theorem S2 in Section S5.3 of the supplementary material). Therefore, the regime $\tilde{d}^{*}\in\tilde{\mathcal{D}}^{*}$ belongs to the solution set $\mathcal{D}^{*}$. Then, since $S^{m}(q,d)$ is a continuous function of $q$, we can consistently estimate the quantile OTRs using Algorithm 1, substituting the estimated optimal decision functions $\hat{f}_{q}$ and the estimated optimal survival functions $\widehat{S}_{h_{n}}(q,\hat{f}_{q})$ with their smoothed counterparts, $\hat{f}_{q}^{m}$ and $\widehat{S}_{h_{n}}^{m}(q,\hat{f}_{q}^{m})$, obtained through survival smoothing.
@@
where $\psi^{m}(a,q;{\pi}^{*},{g}^{*})$ is the smoothed counterpart of $\psi(a,q;{\pi}^{*},{g}^{*})$. The explicit expression for $\psi^{m}(a,q;{\pi}^{*},{g}^{*})$ is given in (3) of the supplementary material.
Then, we apply the classification idea to derive $\hat{f}_{q}^{m}$ and subsequently construct $\widehat{S}_{h_{n}}^{m}(q,\hat{f}_{q}^{m})$ following a similar derivation in Section 2.2. We have provided detailed procedures to compute $\hat{f}_{q}^{m}$ and $\widehat{S}_{h_{n}}^{m}(q,\hat{f}_{q}^{m})$ in Section S2 of the supplementary material. Additionally, a detailed algorithm tailored for discrete outcomes is presented in Algorithm S1 of the supplementary material.
A key advantage of focusing on $\tilde{\mathcal{D}}^{*}$ from the solution set $\mathcal{D}^{*}$ is that it can exclude quantile OTRs that are close to the non-optimal regime and therefore improve the performance when estimating quantile OTRs. When the homoscedastic treatment effect assumption (Assumption 9 in Section 3.2) holds, only the effective quantile OTRs remain in the set $\tilde{\mathcal{D}}^{*}$ (see Theorem S3 in Section S5.3 of the supplementary material), which helps solve the ineffectiveness issue.
@@
Remark 2.3.
Alternative smoothing techniques beyond (7) can be employed provided that the resulting smoothed survival function $S^{m}(q,d)$ is continuous, strictly decreasing, and matches the original survival function $S(q,d)$ at all points $q\in\{v_{1},\dots,v_{l}\}$. Let $\tilde{\mathcal{D}}_{1}^{*}$ and $\tilde{\mathcal{D}}_{2}^{*}$ denote the sets of regimes that maximize the smoothed quantile under two distinct smoothing techniques satisfying the above conditions. We can show that $\tilde{\mathcal{D}}_{1}^{*}=\tilde{\mathcal{D}}_{2}^{*}\subseteq\mathcal{D}^{*}$, demonstrating that different smoothing techniques lead to the same quantile OTRs. More details about the alternative smoothing techniques can be found in Section 5.4 of the supplementary material. We also implement one of the alternative smoothing method and provide a simulation comparing it with (7) in Section 3.5 of the supplementary material. The results reveal no notable differences between the two methods. Therefore, we favor the smoothed survival function in (7), since it is simper and computationally easier.
@@ 3.1 Result for continuous outcomes
Assumption 2: The counterfactual survival functions under treatment $1$ and 0, S(q,1)S({q},1) and S(q,0)S({q},0), are continuous functions of $q$.
Assumption 3: The nuisance models $\hat{\pi}(a|\boldsymbol{x}),\hat{g}(q;\boldsymbol{x},a)$, $q\in(l,r)$, as functions of $\boldsymbol{x}$ and $a$ are in the Vapnik–Chervonenkis (VC) classes $\mathcal{F}_{\pi}$ and $\mathcal{F}_{g}$, respectively.
@@
Assumption 5: For each fixed $(\boldsymbol{x},a)$, $\hat{\pi}(a|\boldsymbol{x})$ and $\hat{g}(q;\boldsymbol{x},a)$ converge in probability to $\pi^{0}(a|\boldsymbol{x})$ and $g^{0}(q;\boldsymbol{x},a)$ uniformly in $q\in(l,r)$ with either $\pi^{0}=\pi^{*}$ or $g^{0}=g^{*}$, where the interval $(l,r)$ contains $(l^{1},r^{1})$ as defined in Algorithm 1. Moreover, there exists a constant $\gamma>0$ such that,
@@
These assumptions were used in the related literature when a classification method is employed to estimate the OTRs zhao2015doubly. VC class (Assumption 3) includes regular parametric models (e.g., the conditional Gaussian model and the generalized linear model) and infinite-dimensional classes when certain smoothness or boundedness are satisfied van1996weak. Assumption 4 is required to ensure the estimator is well-defined with nonzero denominators. Assumption 5 implies that the estimated propensity score $\hat{\pi}$ and the estimated conditional survival function $\hat{g}$ converge to fixed functions $\pi^{0}$ and $g^{0}$, and it requires only one of them is correctly specified, i.e., either $\pi^{0}=\pi^{*}$ or $g^{0}=g^{*}$, suggesting double robustness of our method. If we utilize parametric models to estimate $\pi^{*}$ and $g^{*}$, then $\gamma=1/2$ in Assumption 5.
@@
[…]
@@
In the related literature, assumptions similar to Assumptions 6–8 are typically stated at a fixed value $q$ zhao2012estimating; zhou2017residual; shi2020sparse; ma2011asymptotic. We generalize these assumptions to hold over a neighborhood $Q_{\delta^{*}}$. Since $\delta^{*}$ can be very small, our assumptions are only slightly stronger than those in the existing literature. We require such stronger assumptions to accommodate that the final classification to derive the estimated quantile OTR $\hat{d}_{\hat{q}}$ relies on the estimated optimal quantile $\hat{q}$ rather than a fixed $q$.
Assumption 6 is introduced to guarantee the proper behavior of the classification procedure deriving $\hat{f}_{q}$ in Algorithm 1. The geometric noise exponent $\beta$ (Assumption 6) describes how fast the density of the distance $\Delta(\boldsymbol{X},{q}^{*})$ decays near the decision boundary, i.e., the $\boldsymbol{x}$ satisfying $2\eta(\boldsymbol{x},{q}^{*})-1=0$. When the decay is rapid, there are fewer data points near the boundary, making classification easier, and allowing for the selection of a larger geometric noise exponent $\beta$. An example is distinctly separable data, that is, when $|2\eta(\boldsymbol{X},{q}^{*})-1|>\delta_{0}>0$ with probability $1$ for some constant $\delta_{0}$, and $\eta(\boldsymbol{x},{q})$ is continuous, $\beta$ can be arbitrarily large. Assumption 7 is introduced to ensure the evaluation procedure in Algorithm 1 works properly, i.e., to bound the difference between $\widehat{S}_{h_{n}}(q,\hat{f}_{q})$ and $S(q,d_{q}^{*})$. When $\tau(\boldsymbol{X},{q}^{*})$ has a uniformly bounded density function near 0, and $\tau(\boldsymbol{x},{q})$ is continuous, Assumption 7 holds with $\eta=1$. If there exists some $\delta_{0}>0$ such that $|\tau(\boldsymbol{X},{q}^{*})|>\delta_{0}$ with probability $1$, and $\tau(\boldsymbol{x},{q})$ is continuous, then Assumption 7 holds with an arbitrary large $\eta$. In Assumption 8, for a given $q$, when $S(q_{0},d_{q}^{*})$ is differentiable at point $q_{0}={q}^{*}$ with a nonzero derivative, i.e., the density function for $Y(d_{q}^{*})$ at ${q}^{*}$ is nonzero, both upper and lower bounds for $|S(q_{0},d_{q}^{*})-S({q}^{*},d_{q}^{*})|$ in Assumption 8 can be established with a proper choice of $\delta^{*}$. Thus, Assumption 8 imposes a weaker condition than requiring differentiability.
For simplicity in notations, we also assume $\lambda_{n}$, $\sigma_{n}$, and $h_{n}$ are the same across different q{q}.